OpenStudy (anonymous):

I Will Reward A Metal What is the slope of the line that passes through the pair of points (-5/3, -1) and (-2, 9/2)? A. 2/33 B. -2/33 C. -33/2 D. 33/2

OpenStudy (anonymous):

Please HELP

OpenStudy (alexandervonhumboldt2):

use slope formula: \[\frac{ y_1-y_2 }{ x_1-x_2 }\]

OpenStudy (anonymous):

can you work me through the process please

OpenStudy (alexandervonhumboldt2):

ok: \[\frac{ y_1-y_2 }{ x_1-x_2 }=\frac{ -1-9/2 }{ -5/3+2 }=-\frac{ 11 }{ 2 }*3=-33/2\]

OpenStudy (anonymous):

Thank you so much!!!!1